OpenStudy (anonymous):

GIVING OUT MEDALS AND BECOMING A FAN Find the exact solution to the equation 6-9 log(x+10)=5 x = -2 x = -1 x = 1 x = 19

OpenStudy (anonymous):

\[ 6 - 9 \log(x+10) = 5 \implies 9 \log(x+10) = 1 \implies \log(x+10) = \frac {1}{9}. \]Can you finish from here? (By the way, what base?)

OpenStudy (anonymous):

the base is 9 and could you continue, im sorry:))